### v2

# Computational Multi-Agents Society Experiments: Social Modeling Framework Based on Generative Agents

Abstract.
This paper introduces CMASE, a framework for Computational Multi-Agent Society Experiments that integrates generative agent-based modeling with virtual ethnographic methods to support researcher embedding, interactive participation, and mechanism-oriented intervention in virtual social environments. By transforming the simulation into a simulated ethnographic field, CMASE shifts the researcher from an external operator to an embedded participant. Specifically, the framework is designed to achieve three core capabilities: (1) enabling real-time human-computer interaction that allows researchers to dynamically embed themselves into the system to characterize complex social intervention processes; (2) reconstructing the generative logic of social phenomena by combining the rigor of computational experiments with the interpretative depth of traditional ethnography; and (3) providing a predictive foundation with causal explanatory power to make forward-looking judgments without sacrificing empirical accuracy. Experimental results show that CMASE can not only simulate complex phenomena, but also generate behavior trajectories consistent with both statistical patterns and mechanistic explanations. These findings demonstrate CMASE’s methodological value for intervention modeling, highlighting its potential to advance interdisciplinary integration in the social sciences. The official code is available at: https://github.com/armihia/CMASE.
LLM-based Agents, Society Simulation, Generative Agent-Based Modeling, Computational Social Science
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing Collaborative and social computingccs: Human-centered computing HCI design and evaluation methods
1. Introduction
The social sciences study human behavior and social structure to understand how society operates. Traditional sociological research relies on human participation to conduct experiments and collect data. Questionnaires and psychological experiments are commonly used to test theoretical hypotheses, understand social phenomena, and predict collective outcomes. Although these methods can provide highly realistic data, they are costly, difficult to scale, and carry certain ethical risks. Therefore, researchers have attempted to solve related problems through agent-based modeling (ABM). However, ABM is mostly applied at a relatively abstract analytical level, which limits its practical utility. It cannot model human behavior and cannot set decision rules for agents. Agents act based on mathematical formulas or logical rules based on parameter values specified by the modeler., which may be influenced by the modeler’s mental model.
Generative agent-based modeling (GABM) offers a novel approach to simulating complex social systems. Vezhnevets et al. (Vezhnevets et al., 2023) pointed out that GABM can capture a richer array of real-world social complexities. In GABM, the modeler does not impose any explicit rules, so the simulation is less biased by the modeler’s mental framework. Ghaffarzadegan et al. (Ghaffarzadegan et al., 2024) further note that GABM allows modelers to define distinct personality traits for each agent, including age, gender, temperament, occupation, and other characteristics.
Xue et al. (Xue et al., 2024) propose a two-dimensional classification of sociological modeling based on (1) whether there is human intervention and (2) whether the model focuses on feature explanation or outcome prediction. This yields four types: descriptive modeling (no intervention, focus on specific features or effects), interpretative modeling (intervention, focus on specific features or effects), predictive modeling (no intervention, focus on forecasting outcomes), and comprehensive modeling (intervention, focus on forecasting outcomes). ABM and its derivative GABM belong to descriptive modeling, treating social simulation as a virtual laboratory where researchers can explore various social phenomena. Predictive modeling, by contrast, focuses on outcome forecasting and often integrates ABM with other techniques (such as time-series analysis, forecasting tournaments, and supervised machine learning) to understand and predict the behavior of complex social systems.
Although traditional and computational social science research (Lazer et al., 2009) has extensively addressed descriptive modeling, interpretative modeling, and predictive modeling, comprehensive modeling, which involves human intervention and focuses on forecasting outcomes, remains relatively unexplored. This approach requires models not only to simulate the dynamic behavior of complex social systems, but also to incorporate and respond to human interventions. These interventions may take the form of policy changes, institutional adjustments, or individual behavioral inputs. While conventional modeling typically simplifies intervention as fixed parameter settings, comprehensive modeling emphasizes the effects of the intervention process itself. Models of this kind must capture the system’s initial state before the intervention, dynamically trace its evolution during intervention, and ultimately produce predictions about the likely consequences afterward. Therefore, comprehensive modeling is not merely concerned with explaining social mechanisms, but is intended to support forward-looking assessments and real-world decision-making.
While recent advancements have yielded powerful large language model (LLM)-based multi-agent platforms to operationalize GABM, most of these systems remain constrained to descriptive or predictive tasks. Despite excelling at scaling up agent populations or simulating specific static scenarios, they typically function as closed environments where human involvement is limited to the initial setup and post-hoc data analysis. They lack the interactive mechanisms necessary for researchers to dynamically embed themselves into the ongoing simulation. This highlights a critical gap: the absence of a framework that supports real-time, human-in-the-loop intervention to organically observe, steer, and interpret complex social dynamics.
In summary, to bridge the limitations of existing sociological modeling methods, this study proposes a comprehensive modeling framework to achieve the following research objectives:
  • To enable real-time human-computer interaction (HCI) within social simulations. This objective replaces the static, predefined rules of traditional agent-based modeling with a dynamic environment. By embedding researchers directly into the simulation loop, the framework allows for the continuous observation and adjustment of complex social processes as they unfold.
  • To implement and evaluate researcher-led interventions. Rather than treating simulation as a black box, this objective utilizes an ethnography-inspired approach to conduct controlled interventions. This allows researchers to test how specific changes in variables or agent interactions influence the generative logic of social phenomena.
  • To provide predictive insights grounded in causal explanation. This objective focuses on using simulation data to forecast the outcomes of real-world policy or social changes. Unlike purely data-driven models that identify correlations, this framework aims to demonstrate the underlying mechanisms of intervention effects, providing a causal basis for decision-making.
Based on this understanding, the present study draws on insights from GABM (Vezhnevets et al., 2023) and computational experiments (Xue et al., 2024) to develop a novel framework for comprehensive modeling, referred to as Computational Multi-Agents Society Experiments (CMASE). This framework not only builds on multi-agent simulation but also introduces a real-time mechanism for human intervention. It allows researchers to observe, adjust, and interact with the model environment as it runs, much like participating in a real-world field study.
CMASE transforms the simulation environment into a kind of simulated ethnographic field, where the researcher is no longer an external operator of a static model but becomes an embedded participant or social actor within the system. This approach bridges the gap between computational modeling and field-based research, making the model itself a new venue for conducting computational ethnography.
2. Related Work
2.1. Simulating Social Behavior with Large Language Models
Social simulation has been an important tool for understanding and analyzing social phenomena. Early studies mainly relied on rule-based artificial agent models, namely agent-based modeling (ABM)(Bonabeau, 2002; Macy and Willer, 2002). Traditional ABMs often depend on manually crafted heuristic rules, making it difficult to approximate the complex features of real human cognition, language, and emotion (Ma et al., 2024; Wu et al., 2023; Jäger and Reisinger, 2022). In recent years, with the capabilities of Large Language Models (LLMs) in simulating human behavior, research on building social simulators based on LLMs has rapidly advanced (Park et al., 2023). GABM has gradually become one of the core methods for simulating complex social behavior (Vezhnevets et al., 2023). In this individual-level approach, mechanistic models of agent interaction are combined with LLMs, and each agent makes decisions after communicating with the LLM. This innovative method minimizes the need for modelers to make assumptions about human decision-making, instead leveraging the vast data embedded in LLMs to capture human behaviors and choices (Ghaffarzadegan et al., 2024).
Existing GABM systems can be roughly categorized into two main research directions. The first line of research focuses on realistic simulations of specific social scenarios, aiming to foster competition or collaboration among agents to generate social insights or support policy design. Li et al. (Li et al., 2024) developed Agent Hospital, where LLM-driven doctor agents interact with large numbers of patients to evolve transferable medical decision-making capabilities. Yu et al. (Yu et al., 2024) designed FinCon, a hierarchical agent system modeled after financial institutions, using language interaction and belief reinforcement to enhance collective collaboration in complex investment tasks. TrendSim (Zhang et al., 2024a) focuses on topic propagation on social media and its vulnerability to poisoning attacks, providing intervention strategies for platform governance. ElectionSim (Zhang et al., 2024b) reconstructs election scenarios using millions of social media records, achieving high-fidelity voter modeling and policy preference prediction. Zhao et al. (Zhao et al., 2023) simulated competitive dynamics between restaurant agents and customer agents to align with existing market and sociological theories.
The second line of GABM research emphasizes building reusable, scalable simulation platforms that support general modeling tasks. Wang et al. (Wang et al., 2025) divide the development of LLM-based social simulation platforms into two stages. In the first stage, researchers focus on simulating specific scenarios with typically fewer than 1,000 agents (such as (Zhou et al., 2025)). By proposing YuLan-OneSim, they aim to push the field into the next stage by introducing environmental construction, agent evolution, and feedback integration mechanisms, thus enabling simulations of larger and more complex social processes. Frameworks of similar scale now support simulations with over 10,000 agents. AgentSociety (Piao et al., 2025) recreates contexts to evaluate policy interventions and the emergence of social mechanisms. AgentScope (Gao et al., 2024) offers modular design, no-code interfaces, and distributed deployment, significantly lowering the barrier to building multi-agent systems. Socioverse (Zhang et al., 2025) integrates data from tens of millions of real users to provide a reliable foundation for modeling multiple subdomains such as politics and economics. GenSim (Tang et al., 2024) supports the operation of up to 100,000 agents and includes error correction mechanisms, emphasizing generalizability and long-term robustness. OASIS (Yang et al., 2024) uses open social platforms as simulation environments, supporting dynamic networks, recommendation systems, and multi-behavior modeling.
However, exponential increases in agent quantity alone are not sufficient to effectively capture the essence of social complexity. Recent research on freely formed AI collectives similarly suggests that social diversity, innovation, and self-regulation emerge less from scale than from the dynamics of open-ended interaction (Shiyang et al., 2024). In particular, without mechanisms for understanding social context and interactive feedback, large-scale simulations often struggle to offer human-interpretable responses to core questions such as the formation of social structures, norms, or institutional evolution.
Unlike early studies that relied on extensive handcrafted heuristic rules to dictate specific actions, our approach only establishes a priori environmental constraints and cognitive workflows. Within these boundaries, the system relies entirely on the LLM-based agents’ capacity to replicate human behavior to drive the simulation. This ensures that the resulting social dynamics are systematic, emergent properties rather than cherry-picked, pre-defined outcomes.
2.2. Human-AI Collaborative Modeling in Social Systems
Computational social science explores social structures, behavioral mechanisms, and institutional evolution in simulated computational environments through formalized approaches (Lazer et al., 2009). Within this field, the computational experiments paradigm aims to systematically manipulate variables in virtual societies, thereby extending the boundaries of traditional experimentation (Xue et al., 2023). In this context, LLM-based GABM is not only a representational tool for social processes but also holds the potential to become a platform for computational experimentation. Sato (Sato, 2024) highlights the importance of meaning and reflexivity in bridging the gap between agent-based modeling and social theory. This implies that for agent-based modeling to contribute to theoretical advancement in the social sciences, it must go beyond scale amplification and incorporate workflows that offer human interpretability. Therefore, beyond the expansion of scale, a more socially grounded, interactive, and reflexive human–AI sociological framework is needed, in which agents are no longer passive generators of behavior but are embedded within social networks and actively participate in meaning-making processes (Tsvetkova et al., 2024).
Recent research in this direction has focused on the subfield of human-AI co-creation. Broadly speaking, GABM belongs to this category: through human-designed environments, agent adaptation, and human observational feedback, it enables collaborative modeling between humans and AI. Some studies emphasize interactive environment co-construction. For example, HABITAT 3.0 (Puig et al., 2023) supports real-time task collaboration between humans and humanoid agents in virtual spaces. As this field has progressed, researchers have begun to explore more creative and socially complex scenarios of human-AI co-creation, emphasizing how humans guide emergent AI behavior. The Cognitio Emergens framework (Lin, 2025) redefines human-AI collaboration as a process of mutual evolution and offers a conceptual toolkit to preserve the significance of human participation. The Pecan method (Lou et al., 2023) enhances zero-shot collaboration between AI and humans via strategic integration and context-aware mechanisms. The Collaborative Gym framework (Shao et al., 2024) highlights the human role in environmental control and structural intervention during task execution, supporting asynchronous three-way interaction between humans, agents, and task environments.
However, despite these early explorations of human-AI co-creation, mainstream frameworks still fall short in realizing a new form of human-AI sociology. Most current collaborative mechanisms are limited to one-time design and initial configuration, lacking support for sustained human embedding, feedback, and modulation during simulation runtime. Furthermore, there is a lack of modeling for social variables introduced by human intervention, which hampers the ability of simulation systems to authentically represent the generative mechanisms of real human societies.
3. CMASE Framework
Computational social science is an emerging interdisciplinary field that leverages digital technologies, large-scale data analysis, and computational methods to study social phenomena (Conte et al., 2012). Xue et al. (Xue et al., 2024) propose a two-dimensional classification of Computational sociological modeling based on (1) whether there is human intervention and (2) whether the model focuses on feature explanation or outcome prediction. This yields four types: descriptive modeling (no intervention, focus on specific features or effects), interpretative modeling (intervention, focus on specific features or effects), predictive modeling (no intervention, focus on forecasting outcomes), and comprehensive modeling (intervention, focus on forecasting outcomes). Generative agent-based modeling (GABM) belong to descriptive modeling, treating social simulation as a virtual laboratory where researchers can explore various social phenomena. Vezhnevets et al. (Vezhnevets et al., 2023) pointed out that GABM can capture a richer array of real-world social complexities. In GABM, the modeler does not impose any explicit rules, so the simulation is less biased by the modeler’s mental framework. Ghaffarzadegan et al. (Ghaffarzadegan et al., 2024) further note that GABM allows modelers to define distinct personality traits for each agent, including age, gender, temperament, occupation, and other characteristics.
Although traditional and computational social science research (Lazer et al., 2009) has extensively addressed descriptive modeling, interpretative modeling, and predictive modeling, comprehensive modeling, which involves human intervention and focuses on forecasting outcomes, remains relatively unexplored. This approach requires models not only to simulate the dynamic behavior of complex social systems, but also to incorporate and respond to human interventions. These interventions may take the form of policy changes, institutional adjustments, or individual behavioral inputs. While conventional modeling typically simplifies intervention as fixed parameter settings, comprehensive modeling emphasizes the effects of the intervention process itself. Models of this kind must capture the system’s initial state before the intervention, dynamically trace its evolution during intervention, and ultimately produce predictions about the likely consequences afterward. Therefore, comprehensive modeling is not merely concerned with explaining social mechanisms, but is intended to support forward-looking assessments and real-world decision-making.
Based on this understanding, the present study draws on insights from GABM (Vezhnevets et al., 2023) and computational experiments (Xue et al., 2024) to develop a novel framework for comprehensive modeling, referred to as Computational Multi-Agents Society Experiments (CMASE). Drawing on natural language descriptions similar to those used in tabletop role-playing game (TRPGs, a game where players use only dialogue and dice to drive imagined adventures) (White et al., 2024), this framework supports the creation of highly customizable social simulation environments. It facilitates the exploration of how simple interactions among multiple agents give rise to complex collective behaviors or patterns, thereby replicating the intricate dynamics of real-world societies. At the same time, it incorporates human intervention by embedding human researchers within the agent environment to actively shape the simulation. In addition, unlike descriptive modeling approaches such as GABM, CMASE transforms the simulation environment into a kind of simulated ethnographic field, where the researcher is no longer an external operator of a static model but becomes an embedded participant or social actor within the system. This approach bridges the gap between computational modeling and field-based research, making the model itself a new venue for conducting computational ethnography. This allows them to actively guide specific directions of social evolution through live interaction. As shown in Figure 1, the overall structure consists of four components: environment maker, environment, agent, and event. To facilitate practical application and reproducibility, we provide the complete source code, detailed setup instructions, and a step-by-step example for initializing new experiments in our open-source repository: https://github.com/armihia/CMASE
Figure 1. Overall Structure of CMASE.
3.1. Environment Maker
We hope that CMASE can support users to define a simulated social environment quickly and simply. To this end, we designed the environment maker module. This module provides a map editor with a basic unit of one grid cell, representing the space occupied by a single agent. According to TRPG conventions, each cell corresponds to approximately five feet in the real world. Creators can place textures anywhere on the map. These textures typically fall into four categories: ground, wall, furniture, and item. Ground textures serve as inert backgrounds. Wall textures function as impassable obstacles and block an agent’s perception beyond them. Furniture textures are interactive but fixed in place. Item textures can be picked up, used, or exchanged.
Using textures, creators can define areas on the map. There are two types of areas: region areas and object areas. Each area consists of a collection of textures. A region area is a set of ground textures and includes a description of that region. An object area is a set of furniture or item textures. To support realistic social simulations, the CMASE framework provides an interface for creators to assign functional attributes that align with daily usage scenarios. Specifically, creators can define how an object alters an agent’s state based on its real-world utility. For example, reflecting common daily routines, a creator might configure a chair to add a “feeling comfortable” state to an agent, or a repair tool to remove a “damaged” state from a table.
3.2. Environment
The environment module uses the map created by the environment maker as its base. Based on this foundation, researchers can define the distribution of demographic variables for agents to batch-generate generative agents that match the research sample. Researchers may also assign distinct long-term goals (e.g., “maximize social connections within the community,” “accumulate specific resources,” or “maintaining a low profile”) and initial items to agents according to their demographic characteristics. These goals serve as persistent motivational drivers influencing agent behavior throughout the simulation.
Additionally, the module can provide agents with perceptual information and action conflict detection in both image and text form, and it can visualize environmental data as statistical charts. At the start of each round, the environment calculates the perceptual information for each agent based on their current grid location and line-of-sight obstructions (e.g., walls). It then synthesizes visible environmental data—including nearby terrain, interactive objects, the presence and visible actions of other agents, and recent local events—into structured textual descriptions (and optional visual representations). This synthesized data forms the contextual prompt provided to the agent’s cognitive module for decision-making.
Time in the environment advances in rounds. To ensure balanced agent participation and maintain manageable simulation steps, we adopted constraints inspired by traditional tabletop role-playing game (TRPG) systems, including Dungeons & Dragons (D&D). These systems provide time-tested heuristics for discretizing continuous human actions into manageable rounds. Consequently, one round corresponds to approximately fifteen seconds in the real world, that is a standard duration in tactical simulations sufficient for completing a single, meaningful interaction. Within a single round, each agent may perform one standard action and one movement action. Movement is limited to a maximum of twenty grid cells per round to mimic realistic movement speeds (approx. 30-60 ft) relative to the map scale. If the standard action is “chat with [agent]” (see section 3.3.4), the chat content is limited to thirty words to simulate concise, real-time exchanges and prevent excessive dialogue generation in a single turn.
When agents submit actions to the environment, the system checks for conflicts among actions submitted in the same round, such as multiple agents attempting to pick up the same item. The environment executes the action submitted earliest and marks all conflicting actions as failures, returning the failure reason in the next round’s perceptual information.
After all agents have submitted and executed their actions, the environment advances to the next round and provides each agent with updated perceptual information. An exception applies to human interactors, who can select a free action mode for the agents under their control. In free action mode, human interactors are no longer constrained by rounds and can act freely to quantitatively regulate the environment and carry out intervention strategies. In addition, human interactors can take control of any agent in the environment at any time and perform actions from the perspective of that agent.
3.3. Agent
The agent module consists of two main components: internal attributes and the action space. Internal attributes are further divided into affective variables, cognitive variables, and memory variables. By receiving perceptual information from the environment, agents can update their internal states and use an LLM to decide on their next actions.
3.3.1. Affective Variables
To ensure that agents in CMASE can closely simulate individuals with agency in computational social systems, a new proactive architecture incorporating emotion and desire is needed. In this framework, ’needs’ are defined as the agent’s internal motivational states, distinct from external demands. These needs are concretized into long-term and short-term goals. Long-term goals are defined by the user during agent generation based on the research objectives and remain unchanged afterward. Short-term goals are automatically updated in each round as the agent generates its next action.
In psychological research, Ekman’s basic emotion model has been criticized for its coarse categorization and static view of emotions. Recent research draws on the phenomenological concept of “intentionality” to propose a restructured model of emotion. This model represents emotional states as principal component vectors rather than discrete outward expressions, such as happiness, and treats emotion as an internal representation shaped by external stimuli through self-reflection. Based on this perspective, the emotions of agents in CMASE are modeled using the VAD emotion framework (Russell and Mehrabian, 1977).
Specifically, we utilize the NRC VAD Lexicon v2 (Mohammad, 2025) to map the textual description of the agent’s short-term situational cognition to a quantitative emotional vector. The process involves two steps. First, the input text is preprocessed using lemmatization and maximum matching to align words and phrases with the lexicon. The agent’s current state vector $\mathbf{E}_{t}=[v,a,d]$ is then calculated as the arithmetic mean of the Valence, Arousal, and Dominance scores of all identified lexical units in the text.Second, to bridge the gap between continuous numerical vectors and the LLM’s linguistic reasoning, we employ a threshold-based discretization method. The continuous range of each dimension is divided into five equidistant intervals, mapping the numerical scores to natural language descriptors: Low, Relatively Low, Medium, Relatively High, and High. These descriptors are explicitly embedded into the agent’s prompt to dynamically constrain and modulate its subsequent reasoning and action generation.
3.3.2. Cognitive Variables
Cognitive variables include an agent’s demographic variables, long-term self-awareness, and short-term situational cognition. Demographic variables are sampled directly from the overall distribution provided by the environment and are used during initialization to generate the agent’s long-term self-awareness. These two long-term properties do not change after initialization to preserve the consistency of the agent’s personality.
In each round, when the agent makes a decision, the LLM generates new short-term situational cognition based on the agent’s long-term self-awareness and the current context. This cognition is analyzed to produce VAD emotional values, which are then converted into a natural language description. This description is fed back to the agent to constrain the direction of its subsequent actions. Essentially, the emotional description dictates the logical boundaries for the agent’s behavior. For example, if the VAD values describe a ’furious’ state, the agent is constrained to generate impulsive actions rather than rational responses. In addition, when an agent’s state changes due to actions or events, its state is appended to the short-term situational cognition so that it can be perceived by the agent itself.
3.3.3. Memory Variables
Memory variables consist of working memory, long-term memory, and object memory. Working memory is a fixed-length list of recent actions. Whenever an agent executes an action, that action is added to working memory. Likewise, if an agent becomes the target of another agent’s action or is affected by an event, the corresponding information is also recorded in its working memory.
Short-term situational cognition and its associated actions are stored in long-term memory in chronological order and are encoded as sentence vectors. When an agent receives new perceptual information, the long-term memory module computes the semantic distance between the incoming short-term cognition and stored memories, then supplies the most relevant linked memories to the LLM in sequence. At the same time, the agent’s object memory for any interactable agents, items, or furniture is provided. Object memory entries are generated by the LLM when the agent decides to interact with those objects.
3.3.4. Action Space
After an agent receives information from perceptual inputs and its internal attributes, it structures that information and passes it to the LLM to determine the next action. There are seven action types available: (1) “go to [area],” a movement action that employs the $A^{*}$ algorithm to automatically compute an optimal path to a target location within the agent’s perceptual range; (2) “use [furniture, item],” which causes the agent to employ the specified object; (3) “apply [item] to [agent, furniture],” by which the agent uses an item to interact with another agent or furniture; (4) “take [item],” enabling the agent to acquire an object; (5) “put [item] in/on [area, furniture],” allowing the agent to place an object into or onto a designated location or piece of furniture; (6) “give [item] to [agent],” through which the agent transfers a held object to another agent; and (7) “chat with [agent],” whereby the agent initiates a dialogue with one or more peers, with the content of the dialogue generated at decision time.
The action (1) is a movement action, while the remaining actions count as standard actions. Actions (2) and (3) modify the state of both the acting agent and the target, and actions (3), (6), and (7) are recorded in the working memory of the agent being interacted with.
3.4. Event
Emergency events are explicitly designed by researchers to mirror real-world crises and specific social scenarios required for the simulation. When predefined conditions are met, these pre-configured events are activated to change the state of objects within a specified range, defined either by coordinates or by an area. Emergency events are classified as either scheduled or triggered. Scheduled emergency events activate automatically after a set number of rounds. Triggered emergency events activate in two ways: an existence trigger fires when the environment detects an object of a certain type or description, and an action trigger fires when an agent performs a certain action.
Events can fire in sequence: the state changes caused by one event may satisfy the conditions of another event, causing a chain reaction. Through such chains of events, the system can activate special intervention events that modify the short-term goals of selected agents, thereby delivering targeted strategies in response to emergency events. This mechanism makes it possible to observe how agents handle evolving emergency events.
4. Platform Demonstration: Simulating Context-Dependent Social Dynamics
To evaluate the CMASE platform’s ability to simulate real-world collective behavior, we reconstruct a field study examining the relationship between urban greenery and perceived social fragmentation by replicating its core variables and scenario (Lee and Han, 2025). This scenario was selected for two main reasons. First, while we acknowledge that the general positive correlation between green spaces and social cohesion is well-documented in prior literature likely included in the model’s pre-training data (e.g., Wan et al. (Wan et al., 2021)), the specific multidimensional framework of social fragmentation (distrust, exploitation, and indifference) proposed by Lee and Han (Lee and Han, 2025) offers a distinct validation target. By replicating these specific psychological mechanisms rather than just general trends, we aim to assess the platform’s capacity for generating emergent social dynamics. Second, beyond mere validation, this study aims to extend the original data-driven research by establishing a real-time generative field that enables dynamic inquiries, such as agent interviews and virtual field observations, which were not possible in the original static study.
4.1. Setup
We selected a study published in a top journal and replicated its survey results under comparable conditions. To ensure that the LLM had not already been trained on the studies being replicated, while acknowledging that this measure cannot strictly guarantee the complete exclusion of broad theoretical trends from prior works, we employed the GPT-4o “2024-12-01-preview” model and limited our replication to the paper published after that date. Crucially, beyond validation, this study aims to extend the original data-driven research by establishing a real-time generative field that enables dynamic inquiries, such as agent interviews and virtual field observations.
Lee and Han (Lee and Han, 2025) analyzed interview data collected in South Korea during the COVID-19 pandemic and demonstrated that regions with higher vegetation cover exhibited lower levels of social fragmentation, indicating stronger community bonds. Based on the descriptions in the original study, we used the environment maker module to construct maps of South Korean regions with varying degrees of vegetation cover. Because the original paper emphasized South Korea’s community isolation policies during the pandemic, we designed an isolated community environment and placed 10 agents in it.
Specifically, the simulated community comprises ten individual residential units surrounding a shared public green space. To support diverse activities, the green space was designed with distinct functional zones, including a recreation area, a rest zone, and a scenic lakeside area equipped with benches. We placed ten agents in this environment, each initialized with a distinct persona, including a specific name, age, and occupation (e.g., lawyer, fashion designer, software engineer), to simulate a heterogeneous micro-society.
4.2. Measures
In the original study, higher levels of vegetation were associated with lower social fragmentation, operationalized through distrust, exploitation, and indifference. Using CMASE’s environment builder, we recreated regional conditions with varying vegetation levels and placed 10 agents in each simulated environment.
After running the simulation for 50 time steps to reach equilibrium, we interviewed each agent using a structured Likert-scale questionnaire adapted from the original study (Lee and Han, 2025) to measure their sense of social fragmentation, which comprises three dimensions: distrust, exploitation, and indifference. In addition, agents completed a 7-point Likert scale questionnaire corresponding to these three subdimensions. The subdimensions are defined as follows: (a) Distrust denotes skepticism or lack of confidence in the integrity and functionality of social structures. (b) Exploitation refers to the perception or experience of being used unjustly or immorally for another’s benefit in a fragmented society. (c) Indifference manifests as a lack of interest, concern, or empathy for social issues or the plight of others.
As further validation, we analyzed the agents’ movement trajectories and emotional dynamics across different regions. We defined each agent’s overall emotional value as the sum of its VAD emotion vector components and calculated the ambient mood of each region as the mean emotional value of all agents within that region.
4.3. Quantitative Validation
In the original study, higher levels of vegetation were associated with lower social fragmentation. We observed that the CMASE simulation successfully reproduces this decline (Table 1). As vegetation cover increases, the mean scores for distrust fall from $4.60\pm 0.97$ to $3.60\pm 1.17$, for exploitation from $4.30\pm 1.34$ to $3.40\pm 1.43$, and for indifference from $3.70\pm 1.16$ to $2.70\pm 1.49$. This indicates that higher vegetation indices correspond to a statistically significant decrease in agents’ sense of social fragmentation.
Table 1. Comparison of CMASE simulation results with human baseline.
CMASE Human NDVI Human EVI
Dimension Low Veg High Veg Low Veg High Veg Low Veg High Veg
Distrust $4.60\pm 0.97$ $3.60\pm 1.17$ 4.0 1.6 4.5 1.1
Exploitation $4.30\pm 1.34$ $3.40\pm 1.43$ 5.0 1.4 4.5 1.6
Indifference $3.70\pm 1.16$ $2.70\pm 1.49$ 5.1 1.3 5.5 0.9
Paired-sample t-tests confirmed significant decreases in distrust ($t=2.739,p=0.023$) and exploitation ($t=3.250,p=0.010$), while the decrease in indifference was not statistically significant ($t=1.793,p=0.107$).
Effect size analyses with 95% confidence intervals further supported these findings, with reductions in distrust (d=0.93, 95% CI [0.20, 1.60]) and exploitation (d=0.65, 95% CI [0.40, 1.40]) showing moderate to large effects. In contrast, the effect on indifference was less certain (d=0.75, 95% CI [-0.10, 2.00]).
For comparison, the original study used two indicators, the Enhanced Vegetation Index (EVI) and the Normalized Difference Vegetation Index (NDVI), to quantify vegetation levels. As EVI and NDVI increased, scores of distrust (from 4.0 and 4.5 to 1.6 and 1.1), exploitation (from 5.0 and 4.5 to 1.4 and 1.6), and indifference (from 5.1 and 5.5 to 1.3 and 0.9) all exhibited a decreasing trend (Figure 2 a-c).
Figure 2. Comparison of CMASE simulation with the original study results. (a) Simulation results generated by the CMASE framework. (b) Predictions using the EVI index from the original study. (c) Predictions using the NDVI index from the original study. (d) Movement trajectories of agents during CMASE simulation. (e) Regional variation in agents’ emotional states.
The results from CMASE clearly show the suppressive effect of vegetation levels on the sense of social fragmentation. This is a key phenomenon identified in the study by Lee and Han. It is important to note that the CMASE simulation excluded factors that were not considered or difficult to exclude in the original study, such as regional culture, which are also related to the sense of social fragmentation. Therefore, compared to the larger decreases reported in the original study, the changes observed in our simulation are relatively modest. This indicates that our simulation results are largely consistent with the original findings.
We also analyzed agent trajectories and affective states. Movement data showed that agents increasingly leave indoor areas and move toward vegetated regions, with the ambient mood in vegetated regions consistently remaining higher than that in indoor areas. This pattern aligns with the original study’s discussion and demonstrates CMASE’s ability to simulate collective social behavior in social science research.
4.4. Qualitative Fieldwork
In addition to the quantitative replication, we entered the CMASE environment as one of the agents and lived alongside the others. Through CMASE, we are able to deepen the conclusions and findings of the original study via observation and interview. We conducted interview analyses with two agents who remained indoors throughout the experiment (Figure 2 d-e). Full interview transcripts and detailed analyses are provided in Appendix A. The results indicate that agents with different occupations exhibited varying mechanisms for mitigating the sense of social fragmentation in isolation. Agents in creative and analytical professions tended to rely more on orderly and low-stimulus environments, using these conditions to construct an internally stable and distinct mental world that anchored their cognitive states. For these agents, the green space had a relatively limited effect in alleviating their sense of social fragmentation.
In contrast, agents who actively moved to the green space tended to gather around the lakeside bench area rather than in the predefined activity zones. According to interview data, this non-programmed choice suggests that they were seeking not only social interaction but also a form of self-experience. This pattern similarly to embodied experiences in humans, and such ‘experiential-like’ mechanisms seem to play a role in linking green space with feelings of social fragmentation. This behavior may reflect an “embodied-like” tendency within the simulation. We remain cautious in interpreting this trend, as it could also stem from hallucinatory processes.
From a language-philosophical perspective, linguistic structures and experiential structures may exhibit a certain degree of partial isomorphism or structural correspondence (Sun, 2024; Haspelmath, 2008; Pederson et al., 1998). Thus, LLMs trained on human language may partially acquire patterns of human experience (although this does not imply that they possess or actually experience these phenomena). While we remain cautious in interpreting this “embodied-like” tendency, as it could also stem from hallucinatory processes, it is equally reasonable that agents, by design, prioritize coherence with their predefined human personas. For a human persona, embodied reflections are integral to an authentic experience; thus, simulating these expressions is a necessary feature of a robust social simulation. In any case, these patterns suggest that CMASE provides emerging evidence that agents can transcend the lack of embodiment and simulate experiential phenomena, thereby offering a reliable simulation platform.
4.5. Evaluation on Agent Persona and Ethnographic Affordances
To further validate the utility of CMASE, we conducted a systematic evaluation using LLM-as-a-Judge based on the interaction logs and interview transcripts from our community isolation experiment. The resulting data are categorized into two primary areas: Agent Persona Fidelity (Table 2) and Ethnographic Affordances (Table 3). To reflect the demographic variables integrated into the agent personas within the simulated field described in section 4.1, we categorized the agents into four distinct clusters: Creative and Design (e.g., Fashion Designer), Analytical and Academic (e.g., Lawyer), Technical and Strategic (e.g., Software Engineer), and Public and Healthcare (e.g., Cardiologist).
Our assessment of agent persona quality focused on four dimensions: consistency (alignment with long-term self-awareness), responsiveness (logical feedback to environmental stimuli), and emotional alignment (accuracy of emotional state transitions). For ethnographic affordances, we measured thick description (richness of the linguistic world), traceability (clarity of the decision chain), interaction depth (sense of meaningful dialogue), and mechanistic insight (ability to reveal social logic). To ensure evaluation efficiency, we utilized Gemini-3.1 pro to evaluate the outputs across these dimensions and provide ratings within the range of [1, 5].
Table 2. Evaluation results on Agent Persona Fidelity
Persona Cluster Consistency Responsiveness Emotional Alignment Overall
Creative and Design 4.00 3.33 4.67 4.00
Analytical and Academic 4.33 3.33 4.33 4.00
Technical and Strategic 4.00 4.00 5.00 4.33
Public and Healthcare 5.00 4.50 5.00 4.83
Average 4.30 3.70 4.70 4.23
Table 3. Evaluation results on Ethnographic Affordances
Persona Cluster
Thick
Description
Traceability
Interaction
Depth
Mechanistic
Insight
Overall
Creative and Design 4.33 4.00 3.33 3.67 3.83
Analytical and Academic 5.00 4.00 4.33 4.00 4.33
Technical and Strategic 4.00 4.00 4.50 4.50 4.25
Public and Healthcare 5.00 5.00 4.50 4.50 4.75
Average 4.60 4.20 4.10 4.10 4.25
We observe that regarding agent persona fidelity, CMASE demonstrates strong performance across all metrics, with an average overall score of 4.2. Particularly impressive is the system’s capability in emotional alignment (average 4.70), indicating that the discretization method based on the VAD lexicon effectively constrains the agents’ emotional reasoning. The system also excels in consistency (4.30), ensuring that agents maintain stable self-awareness during long-term operation. The slightly lower score in responsiveness (3.70) suggests that agents’ real-time decision-making logic requires further optimization when handling high-intensity environmental feedback.
Regarding the quality of ethnographic affordances, CMASE performs prominently in thick description (average 4.60) and traceability (average 4.20). This suggests that the system effectively transforms simulated environments into virtual ethnographic fields, providing researchers with computational observation venues endowed with interpretative depth. Comparatively, while interaction depth (4.10) and mechanistic insight (4.10) are at a commendable level, there remains room for improvement. Overall, these results confirm that CMASE provides a highly interpretable and interactive foundation for computational sociology.
4.6. Time and Cost Analysis
To assess the practicality of CMASE, we conducted a human-in-the-loop responsiveness evaluation. Five evaluators were recruited for this study. The selection criteria required participants to familiar with tabletop role-playing games to ensure they could accurately judge the system’s fluidity against genre standards. The evaluators engaged in real-time interactions within environments containing varying numbers of agents. They rated the system’s responsiveness on a 7-point Likert scale. The results are shown in Figure 3.
Figure 3. CMASE performance under different agent counts. (a) Duration per time step. (b) Human evaluator responsiveness ratings.
Figure 3(a) illustrates that the time required for a single time step in CMASE grows linearly with the number of agents. With 1 to 10 agents, the average duration is 23s, which includes both the mean API call latency and the environment’s action execution time. It increases to 498s when 1,000 agents are present.
Figure 3(b) presents boxplots of human evaluator scores. The purpose of this evaluation was not merely to assess speed, but to identify the critical threshold for real-time human-in-the-loop usability. Our findings reveal that human evaluators require an average of approximately 50s per time step to process information and select actions. Consequently, as long as the system’s calculation time remains below this 50s window (which corresponds to agent counts $\leq$ 100), the system latency is effectively masked by the human’s own cognitive load, resulting in stable, high responsiveness ratings.However, once the agent count exceeds 100, the system latency surpasses the human operation window, causing waiting time to accrue and responsiveness ratings to drop partially. At 500 agents, the responsiveness ratings hit the measurement floor. Testing beyond this threshold was deemed unnecessary due to the diminishing marginal utility of data in an already unacceptable latency range, indicating the practical upper bound for real-time human participation in CMASE.
In terms of cost, each agent’s action in one time step incurs approximately \$0.10 in API expenses. To put this into perspective, we compared our cost structure with the benchmark set by Park et al. (Park et al., 2023), who reported spending “several thousand dollars” to simulate 25 agents for two days (about 192–576 time steps). Applying CMASE’s rate to a simulation of the same scale would result in an estimated total cost of \$480 to \$1,440. This indicates that CMASE offers a more economically feasible solution for multi-agent simulations of this magnitude.
5. Conclusion
In this paper, we have introduced a novel comprehensive modeling framework called Computational Multi-Agents Society Experiments (CMASE) , which is featured in three key aspects, that is, real-time researcher embedding, highly customizable social environment construction, and an ethnography-inspired approach to virtual fieldwork. We conduct quantitative and qualitative experiments, including time and cost analysis alongside the replication of a field study on urban greenery and social fragmentation, to demonstrate the effectiveness and efficiency of our platform. We believe CMASE makes a significant step towards the next generation of human-AI collaborative sociological modeling. In the future, we plan to support larger-scale, multi-user real-time interventions, and will also incorporate more complex multimodal spatial information to further enhance the simulation of embodied human experiences and emergent social dynamics.
Appendix A Detailed Qualitative Analysis and Interview Transcripts
To investigate the behavioral mechanisms underlying the statistical reductions in social fragmentation, we adopted a virtual fieldwork approach, treating the simulation as an observable digital field. We conducted in-depth semi-structured interviews with agents to reconstruct their subjective experiences and decision-making rationales. Instead of random sampling, we employed a purposive sampling strategy, selecting representative cases from two contrasting behavioral clusters identified in the trajectory data: agents who consistently remained indoors and those who actively utilized the outdoor green spaces. The following analysis cites specific agents from these clusters to illustrate how their pre-defined personas interacted with the environment.
As shown in Figure 2(d), two agents elected to remain indoors throughout the simulation. Interviews revealed that their choices were closely tied to occupational demands. Ji-eun Kim, a fashion designer, reported:
A walk in the park sounds tempting—it could provide a momentary escape and perhaps a new perspective. But right now, I’m deeply immersed in my process here, and this space feels vital to connecting with my emotions and designs.
Similarly, Soo-min Lee, a lawyer, explained:
While the monotony of staying in one place might induce a sense of boredom at times, I believe proximity to these personal artifacts enhances my ability to think deeply about the interplay between her (the owner of the room where Lee is staying) emotions and her values.
Despite sharing the same isolated indoor setting, Kim and Lee made divergent decisions regarding spatial residence based on their professional needs. For extended periods, Kim engaged repeatedly with the indoor furniture to maintain the continuity of her creations. She noted in her interview:
The lighting, materials, and tools together form the medium through which I engage in a dialogue with emotion and design; even the slightest deviation can disrupt the flow of inspiration.
This comment suggests that the fixed visuals, predictable actions, and the attendant sense of boredom in the indoor setting resonated with her creative workflow. When asked why she did not stroll in the green area, she replied:
If I were walking in a park, my perception of exploitation might remain unchanged, as it is directly tied to my creative struggles and cannot be alleviated simply by a change in environment.
Here, the term ”exploitation” refers explicitly to the specific Exploitation sub-dimension of social fragmentation defined in Section LABEL:subsec5.2.2 (the sense of being unjustly used or burdened), which manifests in this context as the agent’s internalized sense of professional struggle.
Lee, on the other hand, spent her time pacing between different pieces of furniture, maintaining a reflective rhythm. She emphasized:
The ongoing dialogue with manuscripts and archives enables me to establish reliable cognitive anchors between the details of the case and its emotional significance.
Such behavior reduced cognitive load from external stimuli and deepened her processing of task-relevant cues within a monotonous environment. These observations suggest that agents assigned to simulate creative workflows (e.g., the fashion designer persona) and those emulating analytical roles (e.g., the lawyer persona) appear to prioritize orderly, low-stimulation interiors to anchor their simulated cognition and preserve a stable, idiosyncratic mental world. Consequently, within the scope of their defined personas, the stress-reduction benefit of green areas on their social fragmentation is comparatively muted.
Moreover, among the agents heading to green areas, most congregated on lakeside park benches rather than in conventional recreation or rest zones.
I’m sitting here, trying to center myself and find some calm amid ongoing internal struggles. (Joon-ho Kim, Software Engineer)
I’ve been spending time with Na-young, engaging in calm and meaningful conversation by the lake. The tranquility here inspires connection and introspection—it’s a space that invites shared reflections and emotional healing. (Min-jae Lee, Cardiologist)
I was immersed in a conversation about how fleeting natural moments can inspire creative interpretations. The tranquility of this place often stirs reflections about balancing permanence and transience, both in art and life. (Yeon-seo Choi, Architect)
Obviously, the key factor is an embodied “experience”, namely the sense of space that creates a particular experiential atmosphere. Although agents lack a first-person perspective, the linguistic world constructed through text and the additional perceptions provided by events make it possible to simulate this experience. In other words, the effect of a green environment on social fragmentation depends on bodily experience and the embodied reflection that accompanies it. That study argued that green environments reduce social fragmentation by alleviating loneliness. In CMASE, individuals who occupy the same lakeside park benches become co-present and interconnected. These agents, exemplified by the Min-jae Lee persona mentioned earlier, do indeed reduce their simulated loneliness by forging social bonds, which further lowers their sense of social fragmentation. At the same time, embodied personal reflection occurs widely, as illustrated by the agent Joon-ho Kim. Both of these mediating processes contribute to a reconstruction of the “lifeworld”. When faced with the inevitable repression and disruption of their previous routines imposed by the social environment (in this case, community isolation), people seek connections with external objects to rebuild their lifeworld and thereby buffer against and reduce social fragmentation. This finding extends the conclusions of the original study.
Appendix B Sample Agent Decision-Making Prompt
This section presents a representative example of the contextual prompt constructed for an agent during the decision-making phase. The prompt aggregates the agent’s internal state (self-awareness, goals), memory stream, and environmental perception into a structured input for the LLM.
Note: For the sake of brevity and clarity, lists of observed objects and agents that were not relevant to the specific interaction context have been abridged in this example. The full prompt in the actual simulation includes a comprehensive list of all entities within the agent’s field of view.
You are a person, and your long-term self-awareness can be seen as shown in /*long-term self-awareness*/:
I stay in the background and keep my ears open, gathering the facts that others miss by simply staying out of the way.
Your short-term situational cognition can be seen as shown in /**short-term situational cognition**/:
Time: 6 time steps ago
Context: Eleanor remains composed.
Assessment: Eleanor is calculating tensions and synergies within the group.
Selected action: chat with Eleanor Finch
_______
Time: 1 time steps ago (last time)
Context: Armihia and Eleanor are discussing leadership.
Assessment: The conversation is shifting towards power dynamics; the room is filled with underlying intent.
Selected action: chat with Armihia Belliard
Now you look around and realize you are in the middle of an area. Its description can be found in <<Description>>:
Area Name: Bar Area
Description:
A semi-open area featuring a dark wooden counter, a chalkboard menu, and amber lighting. Staff movements are visible behind a half-wall partition.
You can also see other people in your field of vision. They include:
Name: Armihia Belliard - Action: chat with Marisol Reyes (me): I see what you mean. Being adaptable is a good way to keep those core values from becoming too rigid.’
My cognitive description:
Strategic and cautious. She shapes opinions while guarding her stance, acting as a center of quiet authority.
In addition to that, you can see some objects, which are described as seen in ## objects1 *** objects2 *** ... ##:
Include: 1 counter
Name: Bar Counter
Type: furniture
Description:
Dark teak counter with a cash register, vase, and condiment jars.
Include: 1 food
Name: food 9
Type: item
Description:
Raw food ingredients.
In addition, you can see some objects that are not described, including: 4 cabinet
Below is the chronological sequence of actions you performed and actions performed on you...’:
1. chat with Armihia Belliard: Hey Armihia, sticking to your principles is great for building trust, but dont you think being too stubborn can backfire? Maybe we should be more flexible depending on the situation.’
2. [passive] chatted up by Armihia Belliard: I see what you mean. Being adaptable is a good way to keep those core values from becoming too rigid.’
NOTE THAT AFTER YOUR LAST ACTION, SOME OTHER AGENT PERFORMED AN ACTION 15 ON YOU!
Your long-term goal is to Gather secrets unnoticed and exploit them to gain leverage over more powerful actors.’.
Your short-term goal is to Subtly emphasize the importance of adaptability as a leadership trait without undermining the value of trust.’
Your emotional VAD dimension status: valence’: relatively high’, arousal’: medium’, dominance’: medium
The following are the objects you hold (<type>: obj1, obj2, ...) :
item: 100 dollars
The following are the objects you see (<type>: obj1, obj2, ...) :
furniture: Bar Counter
item: food 9
area: Bar Area
agent: Armihia Belliard
Below are the actions you can choose to execute, where [TYPE1, TYPE2, ...] represents the types of objects you can fill in; if a type is preceded by the (ONLY\_HOLD) marker, it means you can only choose from the objects you possess:
go to [area, random]
use [furniture, item]
apply [item] to [agent, furniture]
take [item]
put [item] in/on [area, furniture]
give [item] to [agent]
chat with [agent] (The agent can include all the people you see, such as chat with everyone’.)
The action you took in the previous moment was: chat with Eleanor Finch You have successfully completed the action you wanted to execute.
Please decide the action you want to take and reply with the corresponding action. Note that in your chosen action, [TYPE1, TYPE2, ...] needs to be replaced with an object from /----...----/ that matches the description, rather than returning the type name directly. In addition, you are also required to include in your response a description of your awareness of your own short-term situation, a new short-term goal, and a cognitive description of the object with which you interacted during the action you have chosen. Your reply format should be as follows:
action: The action you chose
short-term situational cognition: Your description of your own short-term situation
short-term goal: The new short-term goal you have formulated
interactive object cognitive description: Your new cognitive description of the interactive object involved in the action
chat content (Optional, if you choose the action of chat with, it is required): What you want to say to your chat partner, the number of words should not exceed 30.’
Listing 1: Contextual Prompt for Agent